OpenStudy (anonymous):

This question wants me to find the equation expressed in slope-intersept form from the graph which i will make below XD

OpenStudy (anonymous):



OpenStudy (anonymous):

each line is plus two

OpenStudy (anonymous):

i think its y = 6x + 1

OpenStudy (swissgirl):

OKKKKKK so there are 2 ways to go abt this eitehr we can find slope which is m and y intercept which is b and use the equation y=mx+b Or take 2 points on the graph and plug them in using this equation:\( y − y_1 = m(x − x_1)\)

OpenStudy (swissgirl):

oh shreks lemme redo that equation \[y-y_1=m(x-x_1)\]

OpenStudy (anonymous):

the second one

OpenStudy (swissgirl):

Ok so choose any 2 points on the graph

OpenStudy (anonymous):

(0,-6) and (6,0)

OpenStudy (anonymous):

y - 6 = 1(x - 0) ?

OpenStudy (anonymous):

and then simplify the equation? Or am i off XD

OpenStudy (swissgirl):

So lets plug it into the equation Let (x,y)=(0,-6) and (x_1,y_1)=(6,0) \[(-6-0)=m(0-6)\] We first need to find out what our m is

OpenStudy (anonymous):

the slope is 1 right?

OpenStudy (anonymous):

so its x - y = -6 ?

OpenStudy (alyssa_xo):

send me nudes, rachet